A particle of charge q and mass m starts from the origin under an electric field E along x axis and magnetic field B along x axis with velocity v. 
After what time the velocity of the particle will be 2v?

Grade:12th Pass

1 Answers

mahima pareek
5 years ago
 
The magnetic field here is parallel to the velocity and hence no effect on it .
By.       F(B) = q [ v ⅹ B]
  So Electric force will affect its velocity
    F(E) = qE 
So acceleration = qE/m
By first equation of motion
i. e .   v = u +a*t 
       2v =  v  + qE/m*t
t =  mv/qE is the answer 
Approve if correct